# Determination of the density in the linear elastic wave equation

### v2

@@ 3. P-waves and related results
with $\nabla\varphi(x_{0})=\xi_{0}$. In order for (7) to hold, we can take
@@
where $A_{P}$ is a scalar function. Under the ray coordinates associated with $g_{P}$,
@@
To determine $A_{P}$, we calculate
@@
Equating the coefficients of order $\mathcal{O}(\varrho)$ in the equation (6) with $i=3$, one obtains
@@
$A_{P}=cJ^{-1/2}c_{P}^{1/2}(\lambda+2\mu)^{-1/2}=cJ^{-1/2}c_{P}^{-1/2}\rho^{-1/2}$
with an arbitrary constant $c$. Since we need to use nontrivial solution, for simplicity, we just take $c=1$.
Next, we collect the coefficients of order $\mathcal{O}(\varrho)$ in the equation (6) with $i=\alpha$. For this, we calculate
@@
$=$ $\partial_{\alpha}(\lambda c_{P}A_{P}c_{P}^{-2})+\lambda c_{P}A_{P}c_{P}^{-2}(\partial_{\gamma}g_{\alpha\beta})g^{\beta\gamma},$
@@
$=$ $\Gamma^{\beta}_{33}\lambda c_{P}^{-1}A_{P}g_{\alpha\beta}c_{P}^{-2}+\Gamma^{\beta}_{\gamma\delta}\lambda c_{P}^{-1}A_{P}g_{\alpha\beta}g^{\gamma\delta}.$
@@
(8) $(\partial_{\gamma}g_{\alpha\beta})g^{\beta\gamma}-\Gamma_{\alpha\beta}^{\beta}-\Gamma^{\beta}_{\gamma\delta}g_{\alpha\beta}g^{\gamma\delta}=0,$
@@
$-\mu c_{P}^{-2}a_{1\alpha}+\rho a_{1\alpha}+\mathrm{i}((\lambda+\mu)c_{P}^{-1}\partial_{\alpha}A_{P}+(\partial_{\alpha}\lambda)c_{P}^{-1}A_{P}-c_{P}^{-2}\partial_{\alpha}c_{P}(3\mu+\lambda)A_{P})=0.$
@@
We need to use the $\mathcal{O}(1)$ terms in the equation (6) with $i=3$, that is,
$\begin{split}&2\mathrm{i}\rho\partial_{t}a_{13}\\ =&\mathrm{i}\left(\partial_{3}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell})+\lambda a_{1k}\varphi_{;\ell}g^{k\ell}\partial_{m}g_{3j}g^{jm}+\partial_{m}(\mu a_{13}\varphi_{;j}+\mu a_{1j}\varphi_{;3})g^{jm}\right)\\ &\mathrm{i}\left(\lambda a_{1k;\ell}g^{k\ell}\varphi_{;3}+\mu(a_{13;j}+a_{1j;3})\varphi_{;m}g^{jm}\right)\\ &-\mathrm{i}\Gamma^{n}_{3m}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell}g_{nj}+\mu a_{1n}\varphi_{;j}+\mu a_{1j}\varphi_{;n})g^{jm}\\ &-\mathrm{i}\Gamma^{n}_{jm}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell}g_{n3}+\mu a_{1n}\varphi_{;3}+\mu a_{13}\varphi_{;n})g^{jm}\\ &+\partial_{m}(\lambda a_{0k;\ell}g^{k\ell}g_{3j}+\mu a_{03;j}+\mu a_{0j;3})g^{jm}\\ &-\Gamma_{3m}^{n}(\lambda a_{0k;\ell}g^{k\ell}g_{nj}+\mu a_{0n;j}+\mu a_{0j;n})g^{jm}\\ &-\Gamma_{jm}^{n}(\lambda a_{0k;\ell}g^{k\ell}g_{n3}+\mu a_{0n;3}+\mu a_{3;n})g^{jm}.\end{split}$
@@
$+\frac{1}{2}\lambda c_{P}^{-2}a_{13}\frac{\partial g_{\alpha\beta}}{\partial\tau}g^{\alpha\beta}-\lambda g^{\alpha\beta}\Gamma_{\alpha\beta}^{\gamma}a_{1\gamma},$
@@
$=$ $(c_{P}^{-3}\partial_{3}c_{P}-\frac{1}{2}c_{P}^{-2}\frac{\partial g_{\alpha\beta}}{\partial\tau}g^{\alpha\beta})(\lambda+2\mu)a_{13}+\Gamma^{\alpha}_{\beta\gamma}\mu a_{1\alpha}g^{\beta\gamma}+\Gamma^{\alpha}_{33}\mu a_{1\alpha}c_{P}^{-2}.$
@@
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
@@
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
$=$ $c_{P}^{-2}(\partial_{3}c_{P})(\lambda+2\mu)(\partial_{3}A_{P})+\frac{1}{2}g^{\delta\gamma}\frac{\partial g_{\delta\gamma}}{\partial\tau}\lambda c_{P}^{-1}(\partial_{3}A_{P})+\sqrt{\rho}\times\text{terms that do notdepend on $\rho$},$
@@
$+c_{P}^{-2}(-c_{P}g^{\alpha\beta}(\partial_{\beta}c_{P}))\mu c_{P}\partial_{\alpha}A_{P}+g^{\gamma\delta}\Gamma^{\alpha}_{\gamma\delta}\mu c_{P}\partial_{\alpha}A_{P}+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
@@
$-g^{\alpha\beta}(\partial_{\beta}c_{P})\mu(\partial_{\alpha}A_{P})+g^{\gamma\delta}\Gamma^{\alpha}_{\gamma\delta}\mu c_{P}(\partial_{\alpha}A_{P})+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
@@
$H_{1}=(\lambda+\mu)g^{\alpha\beta}\partial_{\alpha}a_{1\beta}+g^{\alpha\beta}\partial_{\alpha}\mu a_{1\beta}-(\lambda+4\mu)c_{P}^{-2}\Gamma^{\alpha}_{33}a_{1\alpha}-(\lambda+\mu)g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}a_{1\gamma}-\mu c_{P}^{-2}\Gamma^{\alpha}_{33}a_{1\alpha},$
@@
$\begin{split}H_{2}=&(\lambda+2\mu)c_{P}^{-1}\partial^{2}_{33}A_{P}+\partial_{3}(\lambda+2\mu)c_{P}^{-1}\partial_{3}A_{P}-c_{P}^{-2}(\partial_{3}c_{P})(\lambda+2\mu)(\partial_{3}A_{P})+g^{\alpha\beta}\partial_{\beta}(\mu\partial_{\alpha}(c_{P}A_{P}))\\ &+\frac{1}{2}\partial_{3}\lambda c_{P}^{-1}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}A_{P}-2g^{\alpha\beta}\partial_{\alpha}\mu\partial_{\beta}c_{P}A_{P}\\ &-g^{\alpha\beta}\mu\partial_{\beta}c_{P}\partial_{\alpha}A_{P}+\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}(\lambda+2\mu)c_{P}^{-1}(\partial_{3}A_{P})-g^{\gamma\delta}\Gamma^{\alpha}_{\gamma\delta}\mu c_{P}(\partial_{\alpha}A_{P})\\ &+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.\end{split}$
@@
$-(\lambda+\mu)g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}\frac{\mathrm{i}}{(c_{S}^{2}-c_{P}^{2})\sqrt{\rho}}\left((c_{P}^{2}-3c_{S}^{2})(\partial_{\gamma}\log\sqrt{\rho})c_{P}^{1/2}J^{-1/2}\right)$
@@
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
@@
$\begin{split}H_{1}=&-\mathrm{i}\sqrt{\rho}(c_{P}^{2}-3c_{S}^{2})c_{P}^{1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}\\ &+\mathrm{i}\left(c_{P}^{2}-3c_{S}^{2}-\frac{2(c_{P}^{2}-3c_{S}^{2})}{c_{P}^{2}-c_{S}^{2}}\right)\sqrt{\rho}c_{S}^{2}c_{P}^{1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}\log\sqrt{\rho}\\ &+\mathrm{i}g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}(c_{P}^{2}-3c_{S}^{2})c_{P}^{1/2}J^{-1/2}\partial_{\gamma}\log\sqrt{\rho}\\ &-\mathrm{i}\sqrt{\rho}(c_{P}^{2}-3c_{S}^{2})c_{P}^{-1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}c_{P}\\ &+\mathrm{i}\sqrt{\rho}\left(2c_{S}-\frac{4c_{S}(c_{P}^{2}-5c_{S}^{2})}{c_{P}^{2}-c_{S}^{2}}\right)c_{P}^{-1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}c_{S}\\ &+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.\end{split}$
@@
$+\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}(\lambda+2\mu)c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}\left(-\rho^{-1/2}\partial_{3}\log\sqrt{\rho}\right)-g^{\gamma\delta}\Gamma^{\alpha}_{\gamma\delta}\mu c_{P}J^{-1/2}c_{P}^{-1/2}(-\rho^{-1/2}\partial_{\alpha}\log\sqrt{\rho})$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
@@
$-\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}(\lambda+2\mu)c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}\rho^{-1/2}\partial_{3}\log\sqrt{\rho}+g^{\gamma\delta}\Gamma^{\alpha}_{\gamma\delta}\mu c_{P}J^{-1/2}c_{P}^{-1/2}\rho^{-1/2}\partial_{\alpha}\log\sqrt{\rho}$
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
@@
$+g^{\gamma\delta}\Gamma^{\alpha}_{\gamma\delta}c_{S}^{2}c_{P}^{1/2}J^{-1/2}\rho^{1/2}\partial_{\alpha}\log\sqrt{\rho}+c_{P}^{-1/2}c_{P}^{-2}(3c_{P}^{2})\rho^{1/2}J^{-1/2}\partial_{3}c_{P}\partial_{3}\log\sqrt{\rho}$
@@
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
@@
$+g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}(c_{P}^{2}-4c_{S}^{2})c_{P}^{1/2}J^{-1/2}\partial_{\gamma}\log\sqrt{\rho}+J^{-1/2}c_{P}^{1/2}\rho^{1/2}\partial^{2}_{33}\log\sqrt{\rho}$
@@
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
@@
(12) $\nabla_{\alpha}\nabla_{\beta}\log\sqrt{\rho}=\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}+\frac{1}{2}c_{P}^{-2}\frac{g_{\alpha\beta}}{\partial\tau}\partial_{3}\log\sqrt{\rho}-\Gamma^{\gamma}_{\alpha\beta}\partial_{\gamma}\log\sqrt{\rho},$
@@
$=$ $\rho^{1/2}J^{-1/2}c_{P}^{1/2}\Big{(}-(c_{P}^{2}-4c_{S}^{2})g^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}+(c_{P}^{2}-4c_{S}^{2})g^{\alpha\beta}\Gamma^{\gamma}_{\alpha\beta}\partial_{\gamma}\log\sqrt{\rho}$
@@
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
@@
$+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
Inserting $a_{13}:=c_{P}^{1/2}\rho^{-1/2}J^{-1/2}B$ into the equation (10), we get an equation for $B$ along the unit-speed geodesic $\gamma(s)$ in $(\Omega,g_{P})$,
$2\mathrm{i}\frac{\mathrm{d}B}{\mathrm{d}s}=M_{ij}\dot{\gamma}^{i}\dot{\gamma}^{j}+C,$
where $M$ is a second order tensor
and $C$ does not depend on $\rho$. Recall that $g_{P}=c_{P}^{-2}g_{E}$ is the metric induced by the $P$-wavespeed.
If we solve $B$ along a geodesic $\gamma$ connecting two points $a\in\partial\Omega$ to $b\in\partial\Omega$, then
$B(b)-B(a)=\frac{1}{2\mathrm{i}}\int_{\gamma}(M_{ij}(\gamma(s))\dot{\gamma}^{i}(s)\dot{\gamma}^{j}(s)+C)\mathrm{d}s.$
By carefully analyzing the reflection of waves (with possible mode conversions) (see, for example, [28, 8]) at the boundary or using the argument in [20], one can show that the Dirichlet-to-Neumann map $\Lambda_{\rho}$ determines $B$ on $\partial\Omega$. Similar case is analyzed for acoustic waves in [29]. Therefore we can recover the longitudinal geodesic ray transform of $M$
$\int_{\gamma}M_{ij}(\gamma(s))\dot{\gamma}^{i}(s)\dot{\gamma}^{j}(s)\mathrm{d}s.$
We remark here that the same result was obtained in [20, 6], and then the uniqueness of $\rho$ can be proved assuming, for example, $\lambda=2\mu$ only at isolated points. Here one needs to use the injectivity, up to natural obstructions, of the longitudinal geodesic ray transform (cf., for example, [21, 15, 27]).
4. S-waves and proof of the main theorem
In this section, we will use solutions of the form (4) that represent $S$-waves to derive new results. We work under the ray coordinates associated with the metric $g_{S}=c_{S}^{-2}\mathrm{d}s^{2}$. Now we take
$|\nabla\varphi|^{2}=c_{S}^{-2}$
and $\langle\mathbf{a}_{0},\nabla\varphi\rangle=0$ (or equivalently $a_{03}=0$). Then the identity (7) is satisfied.
We calculate
@@ 4. S-waves and proof of the main theorem
$=$ $c_{S}^{-2}\mu\partial_{3}a_{0\alpha}-c_{S}^{-2}\mu g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{0\beta}$
@@
$=$ $g^{33}\Gamma^{\gamma}_{\alpha 3}\mu a_{0\gamma}+g^{\beta\gamma}\Gamma^{3}_{\alpha\beta}\mu a_{0\gamma}=0,$
@@
$=$ $(g^{33}\Gamma^{3}_{33}+g^{\beta\gamma}\Gamma^{3}_{\beta\gamma})\mu a_{0\alpha}$
$=$ $c_{S}^{-3}\partial_{3}c_{S}\mu a_{0\alpha}-\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}^{-2}\mu a_{0\alpha}.$
As for the $P$-waves, we can also take $\mathbf{a}_{j}$ idependent of $t$. Equating the $\mathcal{O}(\varrho)$-terms in the equation (6) with $i=\alpha$, we obtain
$c_{S}^{-2}\left(2\mu\partial_{3}a_{0\alpha}+(\partial_{3}\mu)a_{0\alpha}-\mu g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{0\beta}+\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}^{-2}\mu a_{0\alpha}-c_{S}^{-1}\partial_{3}c_{S}\mu a_{0\alpha}\right)=0.$
@@
(13) $\frac{\partial\eta_{\alpha}}{\partial\tau}-\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\alpha\beta}}{\partial\tau}\eta_{\gamma}+c_{S}^{-1}\partial_{3}c_{S}\eta_{\alpha}=0.$

### v3

@@ Appendix A Proof of lemmas
In order for (8) to hold, the principal amplitude $\mathbf{a}_{0}$ needs to be parallel with $\nabla\varphi$. We take
@@
Under the ray coordinates associated with $g_{P}$,
@@
To determine $A_{P}$, we use the equation (9) with $i=3$. We calculate
@@
Summing up the above calculations, one obtains
@@
$A_{P}=C_{0}J^{-1/2}c_{P}^{1/2}(\lambda+2\mu)^{-1/2}=C_{0}J^{-1/2}c_{P}^{-1/2}\rho^{-1/2}$
with an arbitrary constant $C_{0}$. This completes the proof of the lemma. Since we need to use nontrivial solutions, one can just take $C_{0}=1$. ∎
Proof of Lemma 2.
In order the construct the subprincipal term of the amplitude, we first use equation (9) with $i=\alpha$. For this, we calculate
@@
$=$ $\partial_{\alpha}(\lambda c_{P}A_{P}c_{P}^{-2})+\lambda c_{P}A_{P}c_{P}^{-2}(\partial_{\varepsilon}g_{\alpha\beta})g^{\beta\varepsilon},$
@@
$=$ $\Gamma^{\beta}_{33}\lambda c_{P}^{-1}A_{P}g_{\alpha\beta}c_{P}^{-2}+\Gamma^{\beta}_{\varepsilon\delta}\lambda c_{P}^{-1}A_{P}g_{\alpha\beta}g^{\varepsilon\delta}.$
@@
(13) $(\partial_{\varepsilon}g_{\alpha\beta})g^{\beta\gamma}-\Gamma_{\alpha\beta}^{\beta}-\Gamma^{\beta}_{\varepsilon\delta}g_{\alpha\beta}g^{\varepsilon\delta}=0,$
@@
$\mu c_{P}^{-2}a_{1\alpha}+\rho a_{1\alpha}+\mathrm{i}((\lambda+\mu)c_{P}^{-1}\partial_{\alpha}A_{P}+(\partial_{\alpha}\lambda)c_{P}^{-1}A_{P}-c_{P}^{-2}\partial_{\alpha}c_{P}(3\mu+\lambda)A_{P})=0.$
@@
Now we have $a_{11}$ and $a_{12}$. In order the obtain $a_{13}$, we subsequently use equation (10) with $i=3$, that is,
$2\mathrm{i}\rho\partial_{t}a_{13}$
$=$ $\mathrm{i}\left(\partial_{3}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell})+\lambda a_{1k}\varphi_{;\ell}g^{k\ell}\partial_{m}g_{3j}g^{jm}+\partial_{m}(\mu a_{13}\varphi_{;j}+\mu a_{1j}\varphi_{;3})g^{jm}\right)$
$\mathrm{i}\left(\lambda a_{1k;\ell}g^{k\ell}\varphi_{;3}+\mu(a_{13;j}+a_{1j;3})\varphi_{;m}g^{jm}\right)$
$-\mathrm{i}\Gamma^{n}_{3m}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell}g_{nj}+\mu a_{1n}\varphi_{;j}+\mu a_{1j}\varphi_{;n})g^{jm}$
$-\mathrm{i}\Gamma^{n}_{jm}(\lambda a_{1k}\varphi_{;\ell}g^{k\ell}g_{n3}+\mu a_{1n}\varphi_{;3}+\mu a_{13}\varphi_{;n})g^{jm}$
$+\partial_{m}(\lambda a_{0k;\ell}g^{k\ell}g_{3j}+\mu a_{03;j}+\mu a_{0j;3})g^{jm}$
$-\Gamma_{3m}^{n}(\lambda a_{0k;\ell}g^{k\ell}g_{nj}+\mu a_{0n;j}+\mu a_{0j;n})g^{jm}$
$-\Gamma_{jm}^{n}(\lambda a_{0k;\ell}g^{k\ell}g_{n3}+\mu a_{0n;3}+\mu a_{3;n})g^{jm}.$
@@
$+\frac{1}{2}\lambda c_{P}^{-2}a_{13}\frac{\partial g_{\alpha\beta}}{\partial\tau}g^{\alpha\beta}-\lambda g^{\alpha\beta}\Gamma_{\alpha\beta}^{\varepsilon}a_{1\varepsilon},$
@@
$=$ $(c_{P}^{-3}\partial_{3}c_{P}-\frac{1}{2}c_{P}^{-2}\frac{\partial g_{\alpha\beta}}{\partial\tau}g^{\alpha\beta})(\lambda+2\mu)a_{13}+\Gamma^{\alpha}_{\beta\varepsilon}\mu a_{1\alpha}g^{\beta\varepsilon}+\Gamma^{\alpha}_{33}\mu a_{1\alpha}c_{P}^{-2}.$
@@
[…]
@@
[…]
$=$ $c_{P}^{-2}(\partial_{3}c_{P})(\lambda+2\mu)(\partial_{3}A_{P})+\frac{1}{2}g^{\delta\varepsilon}\frac{\partial g_{\delta\varepsilon}}{\partial\tau}\lambda c_{P}^{-1}(\partial_{3}A_{P})+\sqrt{\rho}\times\text{terms that do not depend on $\rho$},$
@@
$+c_{P}^{-2}(-c_{P}g^{\alpha\beta}(\partial_{\beta}c_{P}))\mu c_{P}\partial_{\alpha}A_{P}+g^{\varepsilon\delta}\Gamma^{\alpha}_{\varepsilon\delta}\mu c_{P}\partial_{\alpha}A_{P}+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}$
@@
$-g^{\alpha\beta}(\partial_{\beta}c_{P})\mu(\partial_{\alpha}A_{P})+g^{\varepsilon\delta}\Gamma^{\alpha}_{\varepsilon\delta}\mu c_{P}(\partial_{\alpha}A_{P})+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.$
@@
$H_{1}=(\lambda+\mu)g^{\alpha\beta}\partial_{\alpha}a_{1\beta}+g^{\alpha\beta}\partial_{\alpha}\mu a_{1\beta}-(\lambda+4\mu)c_{P}^{-2}\Gamma^{\alpha}_{33}a_{1\alpha}-(\lambda+\mu)g^{\alpha\beta}\Gamma^{\varepsilon}_{\alpha\beta}a_{1\varepsilon}-\mu c_{P}^{-2}\Gamma^{\alpha}_{33}a_{1\alpha},$
@@
$H_{2}=$ $(\lambda+2\mu)c_{P}^{-1}\partial^{2}_{33}A_{P}+\partial_{3}(\lambda+2\mu)c_{P}^{-1}\partial_{3}A_{P}-c_{P}^{-2}(\partial_{3}c_{P})(\lambda+2\mu)(\partial_{3}A_{P})+g^{\alpha\beta}\partial_{\beta}(\mu\partial_{\alpha}(c_{P}A_{P}))$
$+\frac{1}{2}\partial_{3}\lambda c_{P}^{-1}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}A_{P}-2g^{\alpha\beta}\partial_{\alpha}\mu\partial_{\beta}c_{P}A_{P}$
$-g^{\alpha\beta}\mu\partial_{\beta}c_{P}\partial_{\alpha}A_{P}+\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}(\lambda+2\mu)c_{P}^{-1}(\partial_{3}A_{P})-g^{\varepsilon\delta}\Gamma^{\alpha}_{\varepsilon\delta}\mu c_{P}(\partial_{\alpha}A_{P})$
[…]
@@
$-(\lambda+\mu)g^{\alpha\beta}\Gamma^{\varepsilon}_{\alpha\beta}\frac{\mathrm{i}}{(c_{S}^{2}-c_{P}^{2})\sqrt{\rho}}\left((c_{P}^{2}-3c_{S}^{2})(\partial_{\varepsilon}\log\sqrt{\rho})c_{P}^{1/2}J^{-1/2}\right)$
@@
[…]
@@
$\begin{split}H_{1}=&-\mathrm{i}\sqrt{\rho}(c_{P}^{2}-3c_{S}^{2})c_{P}^{1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}\\ &+\mathrm{i}\left(c_{P}^{2}-3c_{S}^{2}-\frac{2(c_{P}^{2}-3c_{S}^{2})}{c_{P}^{2}-c_{S}^{2}}\right)\sqrt{\rho}c_{S}^{2}c_{P}^{1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}\log\sqrt{\rho}\\ &+\mathrm{i}g^{\alpha\beta}\Gamma^{\varepsilon}_{\alpha\beta}(c_{P}^{2}-3c_{S}^{2})c_{P}^{1/2}J^{-1/2}\partial_{\varepsilon}\log\sqrt{\rho}\\ &-\mathrm{i}\sqrt{\rho}(c_{P}^{2}-3c_{S}^{2})c_{P}^{-1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}c_{P}\\ &+\mathrm{i}\sqrt{\rho}\left(2c_{S}-\frac{4c_{S}(c_{P}^{2}-5c_{S}^{2})}{c_{P}^{2}-c_{S}^{2}}\right)c_{P}^{-1/2}J^{-1/2}g^{\alpha\beta}\partial_{\alpha}\log\sqrt{\rho}\partial_{\beta}c_{S}\\ &+\sqrt{\rho}\times\text{terms that do not depend on $\rho$}.\end{split}$
@@
$+\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}(\lambda+2\mu)c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}\left(-\rho^{-1/2}\partial_{3}\log\sqrt{\rho}\right)-g^{\varepsilon\delta}\Gamma^{\alpha}_{\varepsilon\delta}\mu c_{P}J^{-1/2}c_{P}^{-1/2}(-\rho^{-1/2}\partial_{\alpha}\log\sqrt{\rho})$
[…]
@@
$-\frac{1}{2}g^{\alpha\beta}\frac{\partial g_{\alpha\beta}}{\partial\tau}(\lambda+2\mu)c_{P}^{-1}J^{-1/2}c_{P}^{-1/2}\rho^{-1/2}\partial_{3}\log\sqrt{\rho}+g^{\varepsilon\delta}\Gamma^{\alpha}_{\varepsilon\delta}\mu c_{P}J^{-1/2}c_{P}^{-1/2}\rho^{-1/2}\partial_{\alpha}\log\sqrt{\rho}$
[…]
@@
$+g^{\varepsilon\delta}\Gamma^{\alpha}_{\varepsilon\delta}c_{S}^{2}c_{P}^{1/2}J^{-1/2}\rho^{1/2}\partial_{\alpha}\log\sqrt{\rho}+c_{P}^{-1/2}c_{P}^{-2}(3c_{P}^{2})\rho^{1/2}J^{-1/2}\partial_{3}c_{P}\partial_{3}\log\sqrt{\rho}$
@@
[…]
@@
$+g^{\alpha\beta}\Gamma^{\varepsilon}_{\alpha\beta}(c_{P}^{2}-4c_{S}^{2})c_{P}^{1/2}J^{-1/2}\partial_{\varepsilon}\log\sqrt{\rho}+J^{-1/2}c_{P}^{1/2}\rho^{1/2}\partial^{2}_{33}\log\sqrt{\rho}$
@@
[…]
@@
(17) $\nabla_{\alpha}\nabla_{\beta}\log\sqrt{\rho}=\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}+\frac{1}{2}c_{P}^{-2}\frac{g_{\alpha\beta}}{\partial\tau}\partial_{3}\log\sqrt{\rho}-\Gamma^{\varepsilon}_{\alpha\beta}\partial_{\varepsilon}\log\sqrt{\rho},$
@@
$=$ $\rho^{1/2}J^{-1/2}c_{P}^{1/2}\Big(-(c_{P}^{2}-4c_{S}^{2})g^{\alpha\beta}\partial_{\alpha}\partial_{\beta}\log\sqrt{\rho}+(c_{P}^{2}-4c_{S}^{2})g^{\alpha\beta}\Gamma^{\varepsilon}_{\alpha\beta}\partial_{\varepsilon}\log\sqrt{\rho}$
@@
[…]
@@
[…]
Inserting the relation $a_{13}=c_{P}^{1/2}\rho^{-1/2}J^{-1/2}B_{P}$ into the equation (15), we prove the lemma.
Proof of Lemma 3.
With the choice of $\varphi=\varphi_{S}$, we calculate
@@
$=$ $c_{S}^{-2}\mu\partial_{3}a_{0\alpha}-c_{S}^{-2}\mu g^{\beta\varepsilon}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}a_{0\beta},$
@@
$=$ $g^{33}\Gamma^{\varepsilon}_{\alpha 3}\mu a_{0\varepsilon}+g^{\beta\varepsilon}\Gamma^{3}_{\alpha\beta}\mu a_{0\varepsilon}=0,$
@@
$=$ $(g^{33}\Gamma^{3}_{33}+g^{\beta\varepsilon}\Gamma^{3}_{\beta\varepsilon})\mu a_{0\alpha}$
$=$ $c_{S}^{-3}\partial_{3}c_{S}\mu a_{0\alpha}-\frac{1}{2}g^{\beta\varepsilon}\frac{\partial g_{\beta\varepsilon}}{\partial\tau}c_{S}^{-2}\mu a_{0\alpha}.$
As for the $P$-waves, we can also take $\mathbf{a}_{j}$ idependent of $t$. Using equation (9) with $i=\alpha$, we obtain
$c_{S}^{-2}\left(2\mu\partial_{3}a_{0\alpha}+(\partial_{3}\mu)a_{0\alpha}-\mu g^{\beta\varepsilon}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}a_{0\beta}+\frac{1}{2}g^{\beta\varepsilon}\frac{\partial g_{\beta\varepsilon}}{\partial\tau}c_{S}^{-2}\mu a_{0\alpha}-c_{S}^{-1}\partial_{3}c_{S}\mu a_{0\alpha}\right)=0.$
@@
(18) $\frac{\partial\eta_{\alpha}}{\partial\tau}-\frac{1}{2}g^{\beta\varepsilon}\frac{\partial g_{\alpha\beta}}{\partial\tau}\eta_{\varepsilon}+c_{S}^{-1}\partial_{3}c_{S}\eta_{\alpha}=0.$